# Workshop on Mechanistic Interpretability

### v4

@@ 5 Mechanistically Study RAG LLMs for Context Attribution
5.1 Locating Relevant Attention Heads and MLPs
To better understand internal mechanisms by which RAG LLMs attribute generated responses to their relevant context sentences, we systematically investigate specific attention heads and MLP layers involved. Our method combines the ARC-JSD metric described previously (§ 4.1) with the Logit Lens (Nostalgebraist2020, ) to precisely quantify contributions from these internal model components.
Following the ARC-JSD framework in the § 4.1, we apply JSD difference at the level of individual attention heads and MLP layers, comparing their outputs between scenarios involving full context and the ablation of the most relevant context sentence using Eq. 1:
@@ 5.1 Locating Relevant Attention Heads and MLPs
5.2 Mechanistic Insights from Located Attention Heads and MLPs
Applying the methodology described in § 5.1, we conducted experiments across three benchmark datasets (see § 4.2) using various LLM scales. Fig. 2(b) presents the distribution and JSD scores of attention heads identified as most relevant for context attribution in Qwen2-1.5B-Instruct on TyDi QA dataset. Our analysis reveals that the top attention heads contributing to context attribution predominantly reside in the higher layers. This observation holds across most datasets, partially corroborating earlier findings by wu2025retrieval , which indicated that retrieval-related attention heads are typically found in the intermediate and higher layers. Notably, our work expands upon NIAH setting explored by wu2025retrieval by mechanistically evaluating attention heads and MLPs relevance through paraphrasing and contextual integration of RAG LLMs. This setting better reflects real-world RAG applications, where models rarely copy text exactly but instead synthesise and rephrase information from retrieved sources. Additional visualisations and distributions for another Qwen2-7B-IT and Gemma2 models across all datasets are provided in Appendix I. Similarly, Fig. 2(b) illustrates that intermediate and higher MLP layers also significantly contribute to context attribution. This pattern remains consistent across different datasets and model scales within the same LLM family. Corresponding detailed findings for Qwen2-7B-IT and Gemma2 models are available in Appendix I.
@@ 6 Verification of JSD-based Mechanistic Study
Having identified critical RAG components using JSD in § 5.1, we now seek to verify these findings from a different analytical angle. Our JSD approach measures a component’s dependence on context through ablation. As a crucial cross-check, we introduce the semantic gain metric, which measures a component’s direct contribution to the correct answer when the full context is present. This allows us to test a key hypothesis: components that are truly important for attribution should not only depend on the context (high JSD) but also actively use it to improve the answer (high semantic gain). Demonstrating a strong correlation between these metrics will robustly verify that the components we located are indeed the primary drivers of context attribution.
@@ 6.1 Semantic Gains of Attention and MLPs for Context Attribution
Apart from locating relevant attention heads and MLPs using JSD-based metric from the § 5.1, we also know that semantic information of context attribution from attentions and MLPs will be added back to the residual stream from each layer based on the autoregressive language model’s architecture from the § 3 (elhage2021mathematical, ; katz2024backward, ). Based on such properties, we can verify whether the JSD-based metric for attention and MLPs location in the § 5.1 works by projecting the residual stream before and after each layer’s attention and MLPs components into the vocabulary space, and calculating the cosine similarity with the generated response $\mathcal{R}$ to further identify which attention and MLP modules provide higher semantic gains.
@@ 6.2 Mutually Verifying JSD-based Mechanistic Study via the Semantic Gains of Attention and MLPs
Based on the Eq. 4 and Eq. 6, we can locate layer-wise attention and MLP components relevant to context attribution from two different perspectives in the § 5.1 and § 6.1. We can evaluate the correlation of both metrics and further verify the effectiveness of our proposed ARC-JSD metric in the § 4.1 and § 5.1.
Table 3 reports significant Spearman $\rho$ values for overlap between top-10 attention/MLP layers ranked by JSD and by semantic gain. This frequent co-occurrence indicates that both metrics track the same retrieval-driven signal that improves next-token prediction. Intuitively, when a layer genuinely draws on a retrieved sentence $c_{i}$ to write answer, ablating $c_{i}$ (i) alters that layer’s token distribution yielding high JSD and (ii) removes the “helpful push” toward the correct token lowering semantic gain. Layers that merely supply generic syntax or parametric knowledge may boost semantic gain without changing under ablation, so their JSD remains low; the strong overall correlation shows that such cases do not dominate, which further verifies effectiveness of ARC-JSD. In addition, ARC-JSD is practical: it requires only forward passes, avoiding cost and saturation issues of gradient-based saliency (qi2024model, ). Unlike KL (undefined with zero-probability bins) or logit-space $\ell_{2}$ distances (scale-dependent) (ferrando2023explaining, ), JSD is finite, symmetric, scale-free, and measured in interpretable bits.
@@ 7 Case Studies of Located Attention Heads and MLPs
Figure 3: The projection of $\mathbf{x}_{i}^{\ell,\text{mid}}$ and $\mathbf{x}_{i}^{\ell,\text{post}}$ via Logit Lens to vocabulary space from layer 20 to layer 27 of Qwen2-1.5B IT in TyDi QA data sample, where the generated response $\mathcal{R}$ is “A mosquito has two wings.” (See Appendix K for all layer projections). Each cell shows the most probable token decoded via Logit Lens. The colour indicates the probability of the decoded token of the corresponding $\mathbf{x}_{i}^{\ell,\text{mid}}$ or $\mathbf{x}_{i}^{\ell,\text{post}}$.
@@
Based on semantic gains analysis from § 6.2, we visualise projection of middle-residual stream $\mathbf{x}_{i}^{\ell,\text{mid}}$ and post-residual stream $\mathbf{x}_{i}^{\ell,\text{post}}$ via Logit Lens to vocabulary space in Fig. 3 and Appendix K. In Fig. 3, Qwen2-1.5B-IT was given data from TyDi QA dev dataset with context about mosquitos introduction from Wikipedia and query “How many wings does a mosquito have?” as input, and it generates responses “A mosquito has two wings.” as output. Based on our ARC-JSD method, we successfully located top-relevant context sentence, i.e., “Mosquitoes have a slender segmented body, a pair of wings, three pairs of long hair-like legs, feathery antennae, and elongated mouthparts”. When we compare heatmap between $\mathbf{x}_{i}^{\ell,\text{post}}$ and $\mathbf{x}_{i}^{\ell,\text{mid}}$ in Fig. 3 from Layer 20 to Layer 27 (See Appendix K for full heatmap), we find that probability of correct token is increased significantly after $\mathbf{x}_{i}^{\ell,\text{post}}$ compared to $\mathbf{x}_{i}^{\ell,\text{mid}}$, such as ‘wings’ in Layer 23, ‘A’, ‘has’, ‘two’ in Layer 26, and ‘mosquito’, ‘two’, ‘A’ in Layer 27, which aligns with our findings that MLP contribute more parametric knowledge for context attribution in higher layers using JSD-based metric from § 5.2. Moreover, we find that several correct tokens are gradually transferred from their Chinese format to English version in Qwen2, such as ‘一只 (A)’, ‘拥有 (has)’ and ‘翅膀 (wings)’, which is reasonable as Chinese is one of main language resources used in Qwen2 pre- and post-training (yang2024qwen2, ). This finding also matches observations (wu2025the, ) that representations tend to be anchored by semantically-equivalent dominant-language tokens in higher layers.
@@
Control RAGs based on ARC-JSD located Attention & MLPs.
Moreover, we conduct an ablation study to compare the JSD difference of responses by masking the top-10 relevant attention heads and randomly-selected 10 attention heads in Table 5. Generally, ablating attention heads located by using JSD-based metric causes larger JSD scores compared to the random attention heads ablation, which further verifies our proposed ARC-JSD can identify context-attribution-related attention heads.
@@
To further evaluate how the located internal components affect RAG behaviours, we conduct extra experiments on Qwen2 7B IT to demonstrate whether ARC-JSD can reduce hallucinations and maintain factual information at the same time. We select 200 data samples from Hotpot QA where Qwen2 7B IT did not provide a truthful response compared to the ground truth. From Table 4, we can use the JSD attribution scores and semantic gain $G$ as a confidence gate to reduce the hallucination rate by $\approx 39\%$ without hurting factual F1 compared to random gating attention and MLPs.  [^4] This experiment mainly evaluates how our ARC-JSD located attention and MLP affect the hallucinations and factual information in the RAG response based on the well-known agreement that attention heads and MLP focus on retrieval behavior and knowledge storage sun2025redeep , respectively..
@@ 8 Discussion
Comparison JSD with KL, Wasserstein, TV and MMD.
In Fig. 4, we conduct an ablation study among different metrics, which shows that JSD works better than other metrics regarding compute-accuracy trade-off. Here, we discuss some possible reasons: KL divergence will explode whenever ablated run assigns $\approx 0$ probability to a token when full run uses (it is common in deep layers of LLMs). The unbounded scale makes it impossible to compare “how much layer 7 changed” to “how much layer 28 changed”; TV distance is bounded but too coarse, which means that two distributions that swap 5% mass on high-entropy tails give the same TV as two distributions that shift 5% mass off the top-1 token, yet the latter wrecks the answer; Wasserstein needs a distance between tokens. There is no canonical ground metric on a 152K vocabulary (Qwen2-7B-Instruct version), and any choice (e.g., edit distance, embedding cosine, etc.) injects an orthogonal modelling assumption and costs $O(V^{3})$ per layer; MMD always requires a kernel and a feature map to measure a Reproducing kernel Hilbert space (RKHS) norm, which is not tied to likelihood or entropy. It also needs a notion of distance between tokens to build the kernel (See Appendix J for details and examples).
What if all JSD scores are very small?
When all scores are very small, it is the attribution. Small everywhere is not an error, and it means that RAG answers from parametric memory or retrieved passages are irrelevant. In those cases, we prefer to return “no evidence passage was used” rather than force-label the least-bad one. Practically, we can flag the answer with “low-evidence” when all sentence-JSD < 0.02 bits ($\approx$ median noise). The benefit than a threshold is that we can distinguish “no context used” from “weak but present context” without having to guess a universal cut-off. We could use that signal to re-query or warn the user, which is in practice a more faithful and safer behaviour than picking the least-small score.
9 Conclusion
We introduce ARC-JSD, an inference-time JSD–based metric that directly attributes RAG responses to their source sentences with no fine-tuning or surrogate models needed. Across diverse QA benchmarks and instruction-tuned LLMs, ARC-JSD outperforms different baselines in attribution accuracy while cutting computational overhead; when paired with the Logit Lens, it even pinpoints the specific attention heads and MLPs driving those attributions, advancing the mechanistic interpretability and transparency of RAG systems.
@@ Acknowledgement
This work is supported by the Gemma 2 Academic Program GCP Credit Award from Google.
@@ Appendix D Details of the Internal Mechanisms of LLMs
We consider the standard autoregressive Transformer architecture used in LLMs, originally introduced by [32] and subsequently analysed in a series of mechanistic studies [13, 8, 12, 7, 20, 21, 44]. Given a prompt of length $T$, the input tokens $(t_{1},\dots,t_{T})$ from the context-query pair, each drawn from a vocabulary $\mathcal{V}$, are mapped to dd‑dimensional embedding vectors $\mathbf{x}^{0}_{i}\in\mathbb{R}^{d}$, where the embedding matrix $W_{E}\in\mathbb{R}^{|\mathcal{V}|\times d}$.
LLMs normally comprise $L$ identical layers. At layer $\ell$, the residual stream $\mathbf{X}^{\ell}=(\mathbf{x}^{\ell}_{1},\dots,\mathbf{x}^{\ell}_{T}),\quad\mathbf{x}^{\ell}_{i}\in\mathbb{R}^{d}$, acts as a common read–write buffer for both the multi–head attention and the MLP block [8]. For each token $i$, the residual update is
@@
Following the key–value interpretation of MLP layers [13, 8], let $W_{\mathrm{in}}^{\ell}\in\mathbb{R}^{d_{\mathrm{m}}\times d}$ and $W_{\mathrm{out}}^{\ell}\in\mathbb{R}^{d\times d_{\mathrm{m}}}$ denote the input and output weights. Given $\mathbf{x}^{\ell-1}_{i}$, the block first produces coefficients

### v5

@@ 5 Mechanistically Study RAG LLMs for Context Attribution
Locating Relevant Attention Heads and MLPs. To better understand internal mechanisms by which RAG LLMs attribute generated responses to their relevant context sentences, we systematically investigate specific attention heads and MLP layers involved. Our method combines the ARC-JSD metric described previously (§ 4.1) with the Logit Lens (nostalgebraist, 2020) to precisely quantify contributions from these internal model components.
Following the ARC-JSD framework in the § 4.1, we apply JSD difference at the level of individual attention heads and MLP layers in Fig. 1, comparing their outputs between scenarios involving full context and the ablation of the most relevant context sentence using Eq. 1:
@@
Mechanistic Insights from Located Attention Heads and MLPs. Applying methodology described above, we conducted experiments across three benchmark datasets (see § 4.2) using various LLM scales. Fig. 2(b) presents distribution and JSD scores of attention heads identified as most relevant for context attribution in Qwen2-1.5B-Instruct on TyDi QA dataset. Our analysis reveals that the top attention heads contributing to context attribution predominantly reside in the higher layers. This observation holds across most datasets, partially corroborating earlier findings by Wu et al. (2025a), which indicated that retrieval-related attention heads are typically found in the intermediate and higher layers. Notably, our work expands upon NIAH setting explored by Wu et al. (2025a) by mechanistically evaluating attention heads and MLPs relevance through paraphrasing and contextual integration of RAG LLMs. This setting better reflects real-world RAG applications, where models rarely copy text exactly but instead synthesise and rephrase information from retrieved sources. Additional visualisations and distributions for another Qwen2-7B-IT and Gemma2 models across all datasets are provided in Appendix K. Similarly, Fig. 2(b) illustrates that intermediate and higher MLP layers also significantly contribute to context attribution. This pattern remains consistent across different datasets and model scales within the same LLM family. Corresponding detailed findings for Qwen2-7B-IT and Gemma2 models are available in Appendix K.
@@ 6 Verification of JSD-based Mechanistic Study
Having identified critical RAG components using JSD in § 5, we now seek to verify these findings from a different analytical angle. Our JSD approach measures a component’s dependence on context through ablation. As a crucial cross-check, we introduce the semantic gain metric, which measures a component’s direct contribution to the correct answer when the full context is present. This allows us to test a key hypothesis: components that are truly important for attribution should not only depend on the context (high JSD) but also actively use it to improve the answer (high semantic gain). Demonstrating a strong correlation between these metrics will robustly verify that the components we located are indeed the primary drivers of context attribution.
@@ 6.1 Semantic Gains of Attention and MLPs for Context Attribution
Apart from locating relevant attention heads and MLPs using JSD-based metric from the § 5, we also know that semantic information of context attribution from attentions and MLPs will be added back to the residual stream from each layer based on the autoregressive language model’s architecture from the § 3 (Elhage et al., 2021; Katz et al., 2024). Based on such properties, we can verify whether the JSD-based metric for attention and MLPs location in the § 5 works by projecting the residual stream before and after each layer’s attention and MLPs components into the vocabulary space, and calculating the cosine similarity with the generated response $\mathcal{R}$ to further identify which attention and MLP modules provide higher semantic gains.
@@ 6.2 Mutually Verifying JSD-based Mechanistic Study via the Semantic Gains of Attention and MLPs
Based on the Eq. 4 and Eq. 6, we can locate layer-wise attention and MLP components relevant to context attribution from two different perspectives in the § 5 and § 6.1. We can evaluate the correlation of both metrics and further verify the effectiveness of our proposed ARC-JSD metric in the § 4.1 and § 5.
@@
Table 3 reports significant Spearman $\rho$ values for overlap between top-10 attention/MLP layers ranked by JSD and semantic gain. This frequent co-occurrence means that both metrics track the same retrieval signal that improves next-token prediction. Intuitively, when a layer genuinely draws on a retrieved sentence $c_{i}$ to write answers, ablating $c_{i}$ (i) alters that layer’s token distribution yielding high JSD and (ii) removes “helpful push” toward correct tokens lowering semantic gain. Layers that merely supply generic syntax or parametric knowledge may boost semantic gain without changing under ablation, so their JSD remains low; strong overall correlation shows that such cases do not dominate, which further verifies effectiveness of ARC-JSD. Moreover, ARC-JSD is practical: it requires only forward passes, avoiding cost and saturation issues of gradient-based saliency (Qi et al., 2024). Unlike KL (undefined with zero-probability bins) or logit-space $\ell_{2}$ distances (scale-dependent) (Ferrando et al., 2023), JSD is finite, symmetric, scale-free, and measured in interpretable bits.
@@ 7 Case Studies of Located Attention Heads and MLPs
Figure 3: The projection of $\mathbf{x}_{i}^{\ell,\text{mid}}$ and $\mathbf{x}_{i}^{\ell,\text{post}}$ via Logit Lens to vocabulary space from layer 20 to layer 27 of Qwen2-1.5B IT in TyDi QA data sample, where the generated response $\mathcal{R}$ is “A mosquito has two wings.” (See Appendix M for all layer projections). Each cell shows the most probable token decoded via Logit Lens. The colour indicates the probability of the decoded token of the corresponding $\mathbf{x}_{i}^{\ell,\text{mid}}$ or $\mathbf{x}_{i}^{\ell,\text{post}}$.
@@
Based on semantic gains analysis from § 6.2, we visualise projection of middle-residual stream $\mathbf{x}_{i}^{\ell,\text{mid}}$ and post-residual stream $\mathbf{x}_{i}^{\ell,\text{post}}$ via Logit Lens to vocabulary space in Fig. 3 and Appendix M. In Fig. 3, Qwen2-1.5B-IT was given data from TyDi QA dev dataset with context about mosquitos introduction from Wikipedia and query “How many wings does a mosquito have?” as input, and it generates responses “A mosquito has two wings.” as output. Based on our ARC-JSD method, we successfully located top-relevant context sentence, i.e., “Mosquitoes have a slender segmented body, a pair of wings, three pairs of long hair-like legs, feathery antennae, and elongated mouthparts”. When we compare heatmap between $\mathbf{x}_{i}^{\ell,\text{post}}$ and $\mathbf{x}_{i}^{\ell,\text{mid}}$ in Fig. 3 from Layer 20 to Layer 27 (See Appendix M for full heatmap), we find that probability of correct token is increased significantly after $\mathbf{x}_{i}^{\ell,\text{post}}$ compared to $\mathbf{x}_{i}^{\ell,\text{mid}}$, such as ‘wings’ in Layer 23, ‘A’, ‘has’, ‘two’ in Layer 26, and ‘mosquito’, ‘two’, ‘A’ in Layer 27, which aligns with our findings that MLP contribute more parametric knowledge for context attribution in higher layers using JSD-based metric from § 5. Moreover, we find that several correct tokens are gradually transferred from their Chinese format to English version in Qwen2, such as ‘一只 (A)’, ‘拥有 (has)’ and ‘翅膀 (wings)’, which is reasonable as Chinese is one of main language resources used in Qwen2 pre- and post-training (Yang et al., 2024). This finding also matches observations (Wu et al., 2025b) that representations tend to be anchored by semantically-equivalent dominant-language tokens in higher layers.
@@
Control RAGs based on ARC-JSD located Attention & MLPs. We conduct an ablation study to compare JSD difference of responses by masking top-10 relevant attention heads and randomly-selected 10 attention heads in Table 5. Generally, ablating attention heads located by JSD metric causes larger JSD scores compared to random attention heads ablation, which further verifies the effectiveness of our proposed ARC-JSD method.
@@
To further evaluate how located internal components affect RAG behaviours, we conduct extra experiments on Qwen2 7B IT to show whether ARC-JSD can reduce hallucinations and maintain factual information unchanged. We select 200 data samples from Hotpot QA where Qwen2 7B IT did not provide a truthful response compared to ground truth. From Table 4, we use JSD attribution scores and semantic gain $G$ as a confidence gate to reduce hallucination rate by $\approx 39\%$ without hurting factual F1 compared to random gating attention and MLPs [^4] This experiment mainly evaluates how our ARC-JSD located attention and MLP affect the hallucinations and factual information in the RAG response based on the well-known agreement that attention heads and MLP focus on retrieval behavior and knowledge storage Sun et al. (2025), respectively..
Figure 4: Causal activation ablation study of Qwen2-1.5B-IT on TyDi QA dataset for percentage of context attribution accuracy decrease by ablating each attention head and MLP layer.
Causal Activation Ablation Study of ARC-JSD. We further conduct a causal activation ablation study by ablating each attention head and MLP layer for Qwen2-1.5B-IT on TyDi QA dataset. We mainly measure the percentage of context attribution accuracy decrease when each component is ablated. Fig. 4 demonstrates that those attention heads or MLP layers with higher JSD score lead to a larger context attribution accuracy drop, which aligns with findings of mechanistic study in Fig. 2.
Token-level JSD distribution over Model’s Response. We conduct a case study to visualise token-level JSD distribution over the model’s response “A mosquito has two wings.”. Fig. 5 (a) clearly demonstrates that the token “two” contributes the highest JSD score change within the whole response. For another example from TyDi QA with context about “Burntisland Shipbuilding Company”, the query is “Who founded the Burntisland Shipbuilding Company?”, and our ARC-JSD method successfully locates the top-related context sentence, i.e., “Brothers Amos and Wilfrid Ayre founded Burntisland Shipbuilding Co. in 1918 as a First World War emergency shipyard. [1]”. We further visualise model’s response “The Burntisland Shipbuilding Company was founded by brothers Amos and Wilfrid Ayre in 1918 during the First World War as an emergency shipyard.” in the Fig. 5 (b), and we can find that tokens “Am”, “Wil”, “Ay” and “re” contribute more JSD scores, which align with correct answer “Amos and Wilfrid Ayre”.
Figure 5: (a) Token-level JSD distribution over model’s response “A mosquito has two wings.”, where token “two” contributes the highest JSD score. (b) another example with a longer response, where tokens “Am”, “Wil”, “Ay” and “re” contribute more JSD scores.
@@ 8 Discussion
Comparison JSD with KL, Wasserstein, TV and MMD. In Fig. 6, we conduct an ablation study among different metrics, which shows that JSD works better than other metrics regarding compute-accuracy trade-off. Here, we discuss some possible reasons: KL divergence will explode whenever ablated run assigns $\approx 0$ probability to a token when full run uses (it is common in deep layers of LLMs). The unbounded scale makes it impossible to compare “how much layer 7 changed” to “how much layer 28 changed”; TV distance is bounded but too coarse, which means that two distributions that swap 5% mass on high-entropy tails give the same TV as two distributions that shift 5% mass off the top-1 token, yet the latter wrecks the answer; Wasserstein needs a distance between tokens. There is no canonical ground metric on a 152K vocabulary (Qwen2-7B-Instruct version), and any choice (e.g., edit distance, embedding cosine, etc.) injects an orthogonal modelling assumption and costs $O(V^{3})$ per layer;
@@
MMD always requires a kernel and a feature map to measure a Reproducing kernel Hilbert space (RKHS) norm, which is not tied to likelihood or entropy. It also needs a notion of distance between tokens to build the kernel (See Appendix L for details and examples).
What if all JSD scores are very small? When all scores are very small, it is the attribution. Small everywhere is not an error, and it means that RAG answers from parametric memory or retrieved passages are irrelevant. In those cases, we prefer to return “no evidence passage was used” rather than force-label the least-bad one. Practically, we can flag the answer with “low-evidence” when all sentence-JSD << 0.02 bits ($\approx$ median noise). The benefit than a threshold is that we can distinguish “no context used” from “weak but present context” without having to guess a universal cut-off. We could use that signal to re-query or warn the user, which is in practice a more faithful and safer behaviour than picking the least-small score.
Figure 7: The compute-accuracy trade-off on MuSiQue for LLaMA-3.1-8B-IT and Qwen3-Next-80B-A3B-IT with GFLOPs $\mathrm{log}_{10}$ scale per sample.
Generalisation of ARC-JSD to diverse and larger LLMs. We evaluate our ARC-JSD using diverse and larger LLMs, i.e., LLaMA-3.1-8B-IT and Qwen3-Next-80B-A3B-IT, for generalisation evaluation. Fig. 7 shows that our ARC-JSD still has a better compute-accuracy trade-off compared to other baselines, which further demonstrates the effectiveness of ARC-JSD to diverse and larger LLMs based on Fig. 7 and Fig. 2.
Comparison between single-hop QA and multi-hop QA under ARC-JSD. We further take an analysis by comparing the single-hop QA example from the TyDi QA dataset and the multiple-hop QA examples from the Hotpot QA dataset and MuSiQue dataset in Appendix I. We can find that ARC-JSD can directly attribute to the most relevant context that leads to the final answer with the highest JSD score. In addition, the sorted JSD ranking list further indicates how our ARC-JSD method takes causal reasoning to approach the final answer when we list the remaining top relevant contexts with their JSD scores. It brings a possibility that our ARC-JSD can be further used to help track the whole model’s reasoning path for multi-hop QA settings, where we also provide more examples on medical and legal domains with higher demands for interpretability and attribution in Appendix J.
@@ Acknowledgement
This work is supported by the Gemma 2 Academic Program GCP Credit Award from Google and Google Cloud.
Ethics statement
The proposed ARC-JSD uses open-source LLMs publicly available on HuggingFace, and all datasets are public question-answering domains and are publicly available, which did not raise any ethical concerns.
Reproducibility statement
The open-source code of the proposed ARC-JSD is accessible from the anonymous repository: https://github.com/ruizheliUOA/ARC_JSD.
@@ Appendix D Details of the Internal Mechanisms of LLMs
We consider the standard autoregressive Transformer architecture used in LLMs, originally introduced by Vaswani et al. (2017) and subsequently analysed in a series of mechanistic studies (Geva et al., 2021; Elhage et al., 2021; Geva et al., 2022; Dai et al., 2022; Meng et al., 2022a; b; Yuksekgonul et al., 2024). Given a prompt of length $T$, the input tokens $(t_{1},\dots,t_{T})$ from the context-query pair, each drawn from a vocabulary $\mathcal{V}$, are mapped to dd‑dimensional embedding vectors $\mathbf{x}^{0}_{i}\in\mathbb{R}^{d}$, where the embedding matrix $W_{E}\in\mathbb{R}^{|\mathcal{V}|\times d}$.
LLMs normally comprise $L$ identical layers. At layer $\ell$, the residual stream $\mathbf{X}^{\ell}=(\mathbf{x}^{\ell}_{1},\dots,\mathbf{x}^{\ell}_{T}),\quad\mathbf{x}^{\ell}_{i}\in\mathbb{R}^{d}$, acts as a common read–write buffer for both the multi–head attention and the MLP block (Elhage et al., 2021). For each token $i$, the residual update is
@@
Following the key–value interpretation of MLP layers (Geva et al., 2021; Elhage et al., 2021), let $W_{\mathrm{in}}^{\ell}\in\mathbb{R}^{d_{\mathrm{m}}\times d}$ and $W_{\mathrm{out}}^{\ell}\in\mathbb{R}^{d\times d_{\mathrm{m}}}$ denote the input and output weights. Given $\mathbf{x}^{\ell-1}_{i}$, the block first produces coefficients